OpenStudy (anonymous):

y=-x^2-6x-10. Can someone help me graph this. So far, I have (-2.-2) and (-1, -5) but I am not having luck finding another point that works on my graph of 8pts.

jimthompson5910 (jim_thompson5910):

you can use any x value you want to generate a paired/corresponding y value

jimthompson5910 (jim_thompson5910):

say you pick x = 0 y=-x^2-6x-10 y=-(0)^2-6(0)-10 ... replace every x with 0 y = -10 so we know the point (0,-10) is on this graph

jimthompson5910 (jim_thompson5910):

or you could pick x = 1 y=-x^2-6x-10 y=-(1)^2-6(1)-10 ... replace every x with 1 y = -17 the point (1,-17) is on this curve

jimthompson5910 (jim_thompson5910):

generate a bunch of points (say 5 or 6) and plot them all on the same xy coordinate axis then draw a curve through all of them. The more points you plot, the more accurate your graph will be